# Extending Graph Condensation to Multi-Label Datasets: A Benchmark Study

### v1

@@ 1 Introduction
Graph-structured data are fundamental to many real-world applications, including social networks, academic citation networks, chemical molecules, protein-protein interaction networks, mapping services, and product recommendation systems Battaglia et al. (2018); Wu et al. (2020); Zhou et al. (2020); Wu et al. (2022). In these graph structures, nodes represent entities (e.g., users in social networks), while edges represent relationships between these entities (e.g., social connections between users)Zhang et al. (2023).
@@
Graph Neural Networks (GNNs) are a class of deep learning models specifically designed to process and learn from graph-structured data. It utilizes the topology of the graph to capture dependencies and learn meaningful nodes, edges, and graph-level representations. By extending deep learning to non-Euclidean domains, GNNs have achieved state-of-the-arts performance in many graph machine learning applications Zhang et al. (2023); Hamilton et al. (2018). For example, in the recommendation system Li et al. (2019); Mao et al. (2021); Wu et al. (2022); Zhang et al. (2019a), drug discoveries Jiang et al. (2021); Duvenaud et al. (2015); Zhang et al. (2021), fraud account detection Dou et al. (2020); Wang et al. (2021) and traffic forecasting Jiang & Luo (2022). GNNs combine the node features and structure of graph as the input and output the representations of graph information Hamilton et al. (2018); Velickovic et al. (2017); Kipf & Welling (2016). Among all the GNNs techniques, the most broadly followed methods is recursive neighborhood aggregation scheme, which aims to get the representation of each node, respectively Xu et al. (2018). By applying different neighborhood aggregation and graph-level pooling methods, GNNs are capable of handling various of tasks like node classification Xiao et al. (2022), link prediction Hasan & Zaki (2011), and graph classification Zhang et al. (2018).
@@
Even though GNNsZhang et al. (2023; 2019b) are widely used for analyzing graphs, training these models on large-scale graphs is still computationally expensive Gao et al. (2018); Bojchevski et al. (2020). Furthermore, nowadays, graph data is becoming gigantic with numerous nodes and edges in a single graph Zhang et al. (2023). For example, the Twitter user graph has 288M monthly active users as of 3/2015 and an estimated average of 208 followers per user for an estimated total of 60B followers (edges) Ching et al. (2015). Such industry graphs can be two orders of magnitude larger, like hundreds of billions or up to one trillion edges. In other words, the complexity and large scale of graph data pose significant challenges for training GNNsJin et al. (2022b). Despite the scalability challenges, for some specific applications that need re-training multiple times models would encounter more hardships rather than computation cost.
@@
To address these large-scale datasets challenges, graph condensation methods recently Xu et al. (2024); Gao et al. (2024) have gained lots of attention. Graph condensation is motivated by condensing a large graph dataset into a smaller informative synthetic graph, thereby enhancing the scalability and efficiency of GNNs Xu et al. (2024); Gao et al. (2024); Hashemi et al. (2024). The goal is to achieve comparable performance in GNNs using this synthetic small graph, instead of relying on the original large-scale graph Loukas (2018); Jin et al. (2022a). By doing so, condensed synthetic graph enables GNNs to maintain strong predictive performance while significantly reducing the computational resources required for tasks such as training and inference. By eliminating redundant information in original dataset, it makes the synthetic graph more manageable within the constraints of limited computation resources, thereby providing better support for graph data mining tasks and applications such as Continual learning Liu et al. (2023), Network Architecture Search (NAS)Gao et al. (2021), etc. Moreover, take node classification task as an example, the node can be well classified is because GNNs have learned to capture the unique pattern of nodes to distinguish them from other nodes in different classes.
@@
Existing graph condensation works Xu et al. (2024); Gao et al. (2024) investigate different condensing strategies including gradient distance matchingJin et al. (2022b); Yang et al. (2023), trajectory matchingZheng et al. (2024); Zhang et al. (2024), kernel ridge regressionXu et al. (2023); Wang et al. (2024), and distribution matchingLiu et al. (2022; 2023). Even though they have achieved great success, the original and synthetic datasets are all single-label ones. In many real-world scenarios, the nodes of graph can be associated with multiple labels in lots of situations Huang & Zhou (2012); Read et al. (2011). For instance, users are allowed to join multiple groups that respectively represent their diverse interests. Therefore, nodes of social networks Shi et al. (2019) are annotated with several tags profiling users’ preferences in different domains. Similarly, a paper in a citation network Akujuobi et al. (2019) may be associated with multiple research topics, and in the protein-protein interaction networks Zeng et al. (2019), it is used to classify protein functions in biology, which are associated with multiple biological processes, like molecular activities or cellular components. These functions are typically inferred based on how proteins interact with one another, with each interaction providing insight into the role of protein within the cell, such as catalyzing reactions, transporting molecules, or transmitting signals. So in the application of graph datasets, multi-label node classification is a fundamental and practical task in graph data miningShi et al. (2019; 2020); Akujuobi et al. (2019); Zhao et al. (2023). In the multi-label graph scenario, each node not only has content (or features) but also associated with multiple class labels. The goal is to assign one or more labels to each node. However, existing graph condensation techniques are designed to address the single-label graph scalability issue, typically condensing the large graph based on class-wise condensing. In this way, it can be hard for us to get the synthetic small graph from the multi-label graph datasets. While directly applying single-label condensation methods to multi-label graphs would result in a loss of critical information, as the class-wise assumption does not hold. This reminds as the open challenge.
@@ 2.1 Graph Neural Networks
Graph Neural Networks (GNNs) are a class of deep learning models specifically designed to process and learn from graph-structured data. GNNs have achieved outstanding performance in many graph machine learning applications Zhang et al. (2023); Hamilton et al. (2018). For instance, in the recommendation system on social network Li et al. (2019); Mao et al. (2021); Wu et al. (2022); Zhang et al. (2019a), drug discoveries from molecule graphs Jiang et al. (2021); Duvenaud et al. (2015); Zhang et al. (2021), fraud account detection on financial graphs Dou et al. (2020); Wang et al. (2021) and traffic forecasting in transportation graph Jiang & Luo (2022).
GNNs combine the node features and structure of graph as the input and output the representations of graph Hamilton et al. (2018); Velickovic et al. (2017); Kipf & Welling (2016). Among all the GNNs techniques, they broadly follow a recursive neighborhood aggregation scheme to get the representation of each node Xu et al. (2018). Some of the typical GNNs models like the Graph Convolutional Network (GCN) Kipf & Welling (2016) operate by aggregating and transforming information from neighboring nodes layer by layer. Variants like Graph Attenstion Nesworks (GAT) Velickovic et al. (2017) use attention mechanisms to prioritize certain neighbors, while GraphSAGE Hamilton et al. (2017) samples neighborhoods to scale to large graphs. By applying different neighborhood aggregation and graph-level pooling methods, GNNs are capable of handling various of tasks Xiao et al. (2022); Hasan & Zaki (2011); Zhang et al. (2018) like node classification, link prediction, and graph classification.
@@ 2.2 Multi-Label Classification on Graphs
Multi-label classification on graphs is a fundamental task in graph machine learning, where nodes in a graph are associated with multiple labels simultaneously. This task is critical in numerous real-world applications, including social networks, bioinformatics, and recommendation systems. For instance, in social networks, users may belong to multiple interest groups, while in bioinformatics, proteins may participate in multiple biological processes such as molecular functions and cellular activities Huang & Zhou (2012); Read et al. (2011); Shi et al. (2019); Akujuobi et al. (2019). In multi-label classification, the goal is to predict a set of labels for each data point in the dataset. The labels can be thought of as binary variables, where a label is either present (1) or absent (0). Recently, there have been several approaches that use neural networks for multi-label classification.
Traditional multi-label classification methods relied on adapting single-label classification techniques to multi-label scenarios. Popular strategies include problem transformation methods (e.g., binary relevance and classifier chains) and algorithm adaptation approaches, which extend existing classifiers to handle multiple labels directly Read et al. (2011); Tsoumakas & Katakis (2008). These approaches use architectures such as recurrent neural networks (RNNs) and convolutional neural networks (CNNs) to learn representations of the input data and make predictions for the labels. However, these approaches were not designed to exploit the graph structure and are therefore suboptimal for graph data.
With the advent of GNNs, researchers began exploring multi-label classification directly on graph-structured data. GNNs excel in capturing the relational and topological information of graphs, making them well-suited for multi-label classification tasks. For example, GraphSAGE Hamilton et al. (2017) and GAT Velickovic et al. (2017) utilize neighborhood aggregation to generate node embeddings, which can be fed into multi-label classifiers. Specialized models like ML-GCN Chen et al. (2019) integrate graph convolutional layers with multi-label prediction heads to improve performance on multi-label datasets.
@@ 2.3 Graph Condensation
Despite the impressive performance of GNN models, training them on large-scale datasets is often challenging due to high computational costs, especially when retraining or dealing with redundant information. To address these issues, graph condensation techniques have been developed. These methods distill a large-scale graph into a smaller yet informative synthetic graph, improving the scalability and efficiency of GNNs Xu et al. (2024); Gao et al. (2024); Hashemi et al. (2024). The condensation process generates a compact representation that preserves essential topological and feature-based properties from the original graph Loukas (2018); Jin et al. (2022a). This allows GNNs to achieve comparable performance using the condensed graph while significantly reducing the computational resources needed for training and inference.
Existing graph condensation works Xu et al. (2024); Gao et al. (2024) investigate different condensing strategies including gradient distance matchingJin et al. (2022b); Yang et al. (2023), trajectory matchingZheng et al. (2024); Zhang et al. (2024), kernel ridge regressionXu et al. (2023); Wang et al. (2024), and distribution matchingLiu et al. (2022; 2023). Condensed graphs facilitate more efficient model execution without sacrificing the quality of learning, making this technique a promising method for real-world GNN deployments. By eliminating redundant information in original dataset, graph condensation makes the synthetic graph more manageable within the constraints of limited computation resources, thereby providing better support for graph data mining tasks and applications such as Continual learning Liu et al. (2023) and Network Architecture Search (NAS)Gao et al. (2021), etc. Moreover, take node classification as an example, the reason a node can be well classified is that GNNs have learned to capture the unique pattern of nodes to distinguish them from other nodes in different classes.
@@ 3 Problem Formulation
In this work, we adapt the graph condensation techniques to a multi-label setting, trying to get the synthetic multi-label graph. First, we introduce the general framework of graph condensation in Figure 1. Generally, graph condensation methods are using different matching strategies to optimize the synthetic graph $\mathcal{S}$ similar to the original graph $\mathcal{G}$. Despite these various strategies, the key point is to keep the comparable performance of GNNs while training on small graph. Therefore, our objective can be formulated as follows Jin et al. (2022b):
@@
As the specific initialization of model parameters $\theta$ can lead to the overfitting problem Jin et al. (2022b), one common solution Wang et al. (2018) is to generate the synthetic data following a distribution of random initializations of $P_{\theta_{0}}$. Furthermore, the reformulated optimization problem as follows:
@@
[…]
@@ 4.2 SGDD: Structure-broadcasting
Instead of treating $A^{\prime}$ and $X^{\prime}$ in the synthetic graph $\mathcal{S}$ as free parameters, recent work Yang et al. (2023) points out that this way ignores the original structure $A$ in $\mathcal{G}$. To address this problem, they broadcast $A$ as supervision for the generation of $A^{\prime}$. By introducing graphon Gao & Caines (2019); Ruiz et al. (2020); Xia et al. (2023) to matching the different shapes between $A$ and $A^{\prime}$, SGDD use random noise $\mathcal{Z}(N^{\prime})\in\mathbb{R}^{N^{\prime}\times N^{\prime}}$ as input coordinates. Through the generative model to synthesize an adjacency matrix $A^{\prime}$ with $N^{\prime}$ nodes, the process can be formulated as:
@@
where $A$ is supervision and $Dis(\cdot)$ is a metric that measure the difference between $A$ and $A^{\prime}$. By adding the corresponding nodes information from $X^{\prime}$ and $Y^{\prime}$ to the inherent relation learning Pfeiffer III et al. (2014); Shalizi & Thomas (2011), the revised generation model is $GEN(\mathcal{Z}(N^{\prime})\oplus X^{\prime}\oplus Y^{\prime};\Phi)$, $\oplus$ denotes the concatenate operation. In general, SGDD concurrently optimizes the parameters $X^{\prime}$ and $A^{\prime}$. While the refinement of $X^{\prime}$ is achieved through a gradient matching strategy, whereas the $A^{\prime}$ is enhanced using the Laplacian energy distribution(LED) matching technique Tang et al. (2022); Das et al. (2016); Gutman & Zhou (2006). During each step, the other component is frozen to ensure effective refinement, and the overall training loss function can be summarized as:
$\mathcal{M}_{SGDD}=\mathcal{M}_{GCond}+\alpha\mathcal{M}_{structure}+\beta{||A||}_{2},$ (10)
where ${||A||}_{2}$ is proposed as a sparsity regularization term, $\alpha$ and $\beta$ are trade-off parameters.
@@ 4.3 GCDM: Distribution Matching
Inspired by the distribution matching strategy in image dataset condensation Zhao & Bilen (2023), GCDMLiu et al. (2022) adapt the receptive field distribution matching method in GNN. In the graph dataset, with the reproducinf kernal Hilbert space $\mathcal{H}$ defined by GNNs and with in each class $c$, GCDM optmize the maximum mean discrepancy (MMD). Let $\Phi_{\theta}$ be an $L-$layer GNN model parameterized by $\theta$, the loss of GCDM is:

### v2

@@ 1 Introduction
Graph-structured data are fundamental to many real-world applications, including social networks, academic citation networks, chemical molecules, protein-protein interaction networks, mapping services, and product recommendation systems battaglia2018relational; wu2020comprehensive; zhou2020graph; wu2022graph. In these graph structures, nodes represent entities (e.g., users in social networks), while edges represent relationships between these entities (e.g., social connections between users)zhang2023survey.
@@
Graph Neural Networks (GNNs) are a class of deep learning models specifically designed to process and learn from graph-structured data. It utilizes the topology of the graph to capture dependencies and learn meaningful nodes, edges, and graph-level representations. By extending deep learning to non-Euclidean domains, GNNs have achieved state-of-the-arts performance in many graph machine learning applications zhang2023survey; hamilton2018. For example, in the recommendation system Li2019DeepGCNsCG; mao2021ultragcn; wu2022graph; zhang2019star, drug discoveries jiang2021could; duvenaud2015convolutional; zhang2021graph, fraud account detection dou2020enhancing; wang2021review and traffic forecasting jiang2022graph. GNNs combine the node features and structure of graph as the input and output the representations of graph information hamilton2018; velickovic2017graph; kipf2016semi. Among all the GNNs techniques, the most broadly followed methods is recursive neighborhood aggregation scheme, which aims to get the representation of each node, respectively xu2018powerful. By applying different neighborhood aggregation and graph-level pooling methods, GNNs are capable of handling various of tasks like node classification xiao2022graph, link prediction hasan2011survey, and graph classification zhang2018end.
@@
Even though GNNszhang2023survey; zhang2019graph are widely used for analyzing graphs, training these models on large-scale graphs is still computationally expensive gao2018large; bojchevski2020scaling. Furthermore, nowadays, graph data is becoming gigantic with numerous nodes and edges in a single graph zhang2023survey. For example, the Twitter user graph has 288M monthly active users as of 3/2015 and an estimated average of 208 followers per user for an estimated total of 60B followers (edges) ching2015one. Such industry graphs can be two orders of magnitude larger, like hundreds of billions or up to one trillion edges. In other words, the complexity and large scale of graph data pose significant challenges for training GNNsGCond. Despite the scalability challenges, for some specific applications that need re-training multiple times models would encounter more hardships rather than computation cost.
@@
To address these large-scale datasets challenges, graph condensation methods recently xu2024survey; gao2024graph have gained lots of attention. Graph condensation is motivated by condensing a large graph dataset into a smaller informative synthetic graph, thereby enhancing the scalability and efficiency of GNNs xu2024survey; gao2024graph; Hashemi2024ACS. The goal is to achieve comparable performance in GNNs using this synthetic small graph, instead of relying on the original large-scale graph Loukas2018GraphRB; Jin2022CondensingGV. By doing so, condensed synthetic graph enables GNNs to maintain strong predictive performance while significantly reducing the computational resources required for tasks such as training and inference. By eliminating redundant information in original dataset, it makes the synthetic graph more manageable within the constraints of limited computation resources, thereby providing better support for graph data mining tasks and applications such as Continual learning liu2023cat, Network Architecture Search (NAS)gao2021graph, etc. Moreover, take node classification task as an example, the node can be well classified is because GNNs have learned to capture the unique pattern of nodes to distinguish them from other nodes in different classes.
@@
Existing graph condensation works xu2024survey; gao2024graph investigate different condensing strategies including gradient distance matchingGCond; SGDD, trajectory matchingzheng2024structure; zhang2024navigating, kernel ridge regressionxu2023kernel; wang2024fast, and distribution matchingGCDM; liu2023cat. Even though they have achieved great success, the original and synthetic datasets are all single-label ones. In many real-world scenarios, the nodes of graph can be associated with multiple labels in lots of situations huang2012multi; read2011classifier. For instance, users are allowed to join multiple groups that respectively represent their diverse interests. Therefore, nodes of social networks shi2019mlne are annotated with several tags profiling users’ preferences in different domains. Similarly, a paper in a citation network akujuobi2019collaborative may be associated with multiple research topics, and in the protein-protein interaction networks zeng2019graphsaint, it is used to classify protein functions in biology, which are associated with multiple biological processes, like molecular activities or cellular components. These functions are typically inferred based on how proteins interact with one another, with each interaction providing insight into the role of protein within the cell, such as catalyzing reactions, transporting molecules, or transmitting signals. So in the application of graph datasets, multi-label node classification is a fundamental and practical task in graph data miningshi2019mlne; shi2020multi; akujuobi2019collaborative; zhao2023multi. In the multi-label graph scenario, each node not only has content (or features) but also associated with multiple class labels. The goal is to assign one or more labels to each node. However, existing graph condensation techniques are designed to address the single-label graph scalability issue, typically condensing the large graph based on class-wise condensing. In this way, it can be hard for us to get the synthetic small graph from the multi-label graph datasets. While directly applying single-label condensation methods to multi-label graphs would result in a loss of critical information, as the class-wise assumption does not hold. This reminds as the open challenge.
@@ 2.1 Graph Neural Networks
Graph Neural Networks (GNNs) are a class of deep learning models specifically designed to process and learn from graph-structured data. GNNs have achieved outstanding performance in many graph machine learning applications zhang2023survey; hamilton2018. For instance, in the recommendation system on social network Li2019DeepGCNsCG; mao2021ultragcn; wu2022graph; zhang2019star, drug discoveries from molecule graphs jiang2021could; duvenaud2015convolutional; zhang2021graph, fraud account detection on financial graphs dou2020enhancing; wang2021review and traffic forecasting in transportation graph jiang2022graph.
GNNs combine the node features and structure of graph as the input and output the representations of graph hamilton2018; velickovic2017graph; kipf2016semi. Among all the GNNs techniques, they broadly follow a recursive neighborhood aggregation scheme to get the representation of each node xu2018powerful. Some of the typical GNNs models like the Graph Convolutional Network (GCN) kipf2016semi operate by aggregating and transforming information from neighboring nodes layer by layer. Variants like Graph Attenstion Nesworks (GAT) velickovic2017graph use attention mechanisms to prioritize certain neighbors, while GraphSAGE hamilton2017inductive samples neighborhoods to scale to large graphs. By applying different neighborhood aggregation and graph-level pooling methods, GNNs are capable of handling various of tasks xiao2022graph; hasan2011survey; zhang2018end like node classification, link prediction, and graph classification.
@@ 2.2 Multi-Label Classification on Graphs
Multi-label classification on graphs is a fundamental task in graph machine learning, where nodes in a graph are associated with multiple labels simultaneously. This task is critical in numerous real-world applications, including social networks, bioinformatics, and recommendation systems. For instance, in social networks, users may belong to multiple interest groups, while in bioinformatics, proteins may participate in multiple biological processes such as molecular functions and cellular activities huang2012multi; read2011classifier; shi2019mlne; akujuobi2019collaborative. In multi-label classification, the goal is to predict a set of labels for each data point in the dataset. The labels can be thought of as binary variables, where a label is either present (1) or absent (0). Recently, there have been several approaches that use neural networks for multi-label classification.
Traditional multi-label classification methods relied on adapting single-label classification techniques to multi-label scenarios. Popular strategies include problem transformation methods (e.g., binary relevance and classifier chains) and algorithm adaptation approaches, which extend existing classifiers to handle multiple labels directly read2011classifier; tsoumakas2008multi. These approaches use architectures such as recurrent neural networks (RNNs) and convolutional neural networks (CNNs) to learn representations of the input data and make predictions for the labels. However, these approaches were not designed to exploit the graph structure and are therefore suboptimal for graph data.
With the advent of GNNs, researchers began exploring multi-label classification directly on graph-structured data. GNNs excel in capturing the relational and topological information of graphs, making them well-suited for multi-label classification tasks. For example, GraphSAGE hamilton2017inductive and GAT velickovic2017graph utilize neighborhood aggregation to generate node embeddings, which can be fed into multi-label classifiers. Specialized models like ML-GCN chen2019multi integrate graph convolutional layers with multi-label prediction heads to improve performance on multi-label datasets.
@@ 2.3 Graph Condensation
Despite the impressive performance of GNN models, training them on large-scale datasets is often challenging due to high computational costs, especially when retraining or dealing with redundant information. To address these issues, graph condensation techniques have been developed. These methods distill a large-scale graph into a smaller yet informative synthetic graph, improving the scalability and efficiency of GNNs xu2024survey; gao2024graph; Hashemi2024ACS. The condensation process generates a compact representation that preserves essential topological and feature-based properties from the original graph Loukas2018GraphRB; Jin2022CondensingGV. This allows GNNs to achieve comparable performance using the condensed graph while significantly reducing the computational resources needed for training and inference.
Existing graph condensation works xu2024survey; gao2024graph investigate different condensing strategies including gradient distance matchingGCond; SGDD, trajectory matchingzheng2024structure; zhang2024navigating, kernel ridge regressionxu2023kernel; wang2024fast, and distribution matchingGCDM; liu2023cat. Condensed graphs facilitate more efficient model execution without sacrificing the quality of learning, making this technique a promising method for real-world GNN deployments. By eliminating redundant information in original dataset, graph condensation makes the synthetic graph more manageable within the constraints of limited computation resources, thereby providing better support for graph data mining tasks and applications such as Continual learning liu2023cat and Network Architecture Search (NAS)gao2021graph, etc. Moreover, take node classification as an example, the reason a node can be well classified is that GNNs have learned to capture the unique pattern of nodes to distinguish them from other nodes in different classes.
@@ 3 Problem Formulation
In this work, we adapt the graph condensation techniques to a multi-label setting, trying to get the synthetic multi-label graph. First, we introduce the general framework of graph condensation in Figure 1. Generally, graph condensation methods are using different matching strategies to optimize the synthetic graph $\mathcal{S}$ similar to the original graph $\mathcal{G}$. Despite these various strategies, the key point is to keep the comparable performance of GNNs while training on small graph. Therefore, our objective can be formulated as follows GCond:
@@
As the specific initialization of model parameters $\theta$ can lead to the overfitting problem GCond, one common solution wang2018dataset is to generate the synthetic data following a distribution of random initializations of $P_{\theta_{0}}$. Furthermore, the reformulated optimization problem as follows:
@@
$\text{s.t.}\quad\theta^{*}=\arg\min_{\theta(\theta_{0})}\mathcal{L}({GNN}_{\theta})].$
@@ 4.2 SGDD: Structure-broadcasting
Instead of treating $A^{\prime}$ and $X^{\prime}$ in the synthetic graph $\mathcal{S}$ as free parameters, recent work SGDD points out that this way ignores the original structure $A$ in $\mathcal{G}$. To address this problem, they broadcast $A$ as supervision for the generation of $A^{\prime}$. By introducing graphon gao2019graphon; ruiz2020graphon; xia2023implicit to matching the different shapes between $A$ and $A^{\prime}$, SGDD use random noise $\mathcal{Z}(N^{\prime})\in\mathbb{R}^{N^{\prime}\times N^{\prime}}$ as input coordinates. Through the generative model to synthesize an adjacency matrix $A^{\prime}$ with $N^{\prime}$ nodes, the process can be formulated as:
@@
where $A$ is supervision and $Dis(\cdot)$ is a metric that measure the difference between $A$ and $A^{\prime}$. By adding the corresponding nodes information from $X^{\prime}$ and $Y^{\prime}$ to the inherent relation learning pfeiffer2014attributed; shalizi2011homophily, the revised generation model is $GEN(\mathcal{Z}(N^{\prime})\oplus X^{\prime}\oplus Y^{\prime};\Phi)$, $\oplus$ denotes the concatenate operation. In general, SGDD concurrently optimizes the parameters $X^{\prime}$ and $A^{\prime}$. While the refinement of $X^{\prime}$ is achieved through a gradient matching strategy, whereas the $A^{\prime}$ is enhanced using the Laplacian energy distribution(LED) matching technique tang2022rethinking; das2016distribution; gutman2006laplacian. During each step, the other component is frozen to ensure effective refinement, and the overall training loss function can be summarized as:
[…]
@@ 4.3 GCDM: Distribution Matching
Inspired by the distribution matching strategy in image dataset condensation zhao2023dataset, GCDMGCDM adapt the receptive field distribution matching method in GNN. In the graph dataset, with the reproducinf kernal Hilbert space $\mathcal{H}$ defined by GNNs and with in each class $c$, GCDM optmize the maximum mean discrepancy (MMD). Let $\Phi_{\theta}$ be an $L$-layer GNN model parameterized by $\theta$, the loss of GCDM is: